# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ V-A2 Multi-Frame Case: Dynamic Channel Evolution
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{RKHS-MF}}&=[\bm{R}_{\underline{zs}}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{R}_{\underline{z}}+\lambda\bm{I}_{L}]^{-1}\\ &=\left(\frac{1}{L}\underline{\bm{S}}\bm{K}+\lambda\bm{W}^{\prime\mathsf{H}}\right)\cdot\left(\frac{1}{L}\bm{K}^{2}+\lambda\bm{I}_{L}\right)^{-1},\end{array}$ (57)
where $\lambda\geq 0$ is a tuning parameter to control the similarity between $\bm{W}$ and $\bm{W}^{\prime}$; cf. Claim 1. normal-$\square$
The remaining distributional robustness modeling and analyses on (56) against the uncertainties in $\hat{\bm{R}}_{\underline{z}}$, $\hat{\bm{R}}_{\underline{xz}}$, and $\hat{\bm{R}}_{\underline{s}}$ are technically straightforward; see Subsection IV-C. Therefore, we omit them here.
@@ VI Experiments
The following eleven methods are implemented in the experiments: 1) Wiener: Wiener beamformer (13), upper expression; 2) Wiener-DL: Wiener beamformer with diagonal loading (27), upper expression; 3) Wiener-DR: Distributionally robust Wiener beamformer (37) and (41); 4) Wiener-CE: Channel-estimation-based Wiener beamformer (13), lower expression; 5) Wiener-CE-DL: Channel-estimation-based Wiener beamformer with diagonal loading (27), lower expression; 6) Wiener-CE-DR: Distributionally robust channel-estimation-based Wiener beamformer (31) and (33); 7) Capon: Capon beamformer (28) for $\epsilon_{1}=0$; 8) Capon-DL: Capon beamformer with diagonal loading (28); 9) ZF: Zero-forcing beamformer where $\bm{W}_{\text{ZF}}\coloneqq(\hat{\bm{H}}^{\mathsf{H}}\hat{\bm{H}})^{-1}\hat{\bm{H}}^{\mathsf{H}}$ and $\hat{\bm{H}}$ denotes the estimated channel matrix; 10) Kernel: Kernel receiver (50) with all $\epsilon$’s in (49) equal to zeroes; and 11) Kernel-DL: Kernel receiver with diagonal loading (54). Note that the diagonal-loading-based methods are particular cases of distributionally robust beamformers; see, e.g., Examples 1 and 6. The deep-learning-based (DL-based) methods in Subsection V-B are not implemented in this section because they have been deeply studied in our previous publications, e.g., [10, 12]; we only comment on the advantages and disadvantages of DL-based methods compared with the listed eleven methods in Section VII (Conclusions).
@@
When covariance matrix $\bm{P}$ of transmitted signal $\mathbf{s}$ is unknown for the receiver (e.g., in ISAC systems, $\bm{P}$ needs to vary from one frame to another for sensing), $\bm{P}$ is estimated by the sample covariance matrix $\hat{\bm{P}}=\bm{S}\bm{S}^{\mathsf{H}}/L$. The channel matrix $\bm{H}$ is estimated using the minimum mean-squared error method, i.e., $\hat{\bm{H}}=\bm{X}\bm{S}^{\mathsf{H}}(\bm{S}\bm{S}^{\mathsf{H}})^{-1}$. Covariance matrix $\bm{R}$ of channel noise $\mathbf{v}$ is estimated using the least-square method, i.e., $\hat{\bm{R}}=(\bm{X}-\hat{\bm{H}}\bm{S})(\bm{X}-\hat{\bm{H}}\bm{S})^{\mathsf{H}}/L$. The matrices $\hat{\bm{P}}$, $\hat{\bm{H}}$, and $\hat{\bm{R}}$ are therefore uncertain compared to their true (but unknown) values $\bm{P}$, $\bm{H}$, and $\bm{R}$, respectively. The matrices $\hat{\bm{P}}$, $\hat{\bm{H}}$, and $\hat{\bm{R}}$ are used in beamformers such as the channel-estimation-based Wiener beamformer (27), the Capon beamformer, and the zero-forcing beamformer.
@@
The beamformers are determined on the training data set (i.e., pilot data). The performance evaluation method of beamformers is mean-squared estimation error (MSE) on the test data set (i.e., non-pilot communication data); when the transmitted discrete symbols are from constellations, the symbol-error rate (SER) is used. As data-driven machine learning methods, all parameters (e.g., uncertainty quantification coefficients $\epsilon$’s) of beamformers can be tuned using the popular cross-validation (e.g., one-shot cross-validation) method. The parameters can also be empirically tuned to save training times because cross-validation imposes a significant computational burden. This paper mainly uses the empirical tuning method (i.e., trial-and-error) to tune each beamformer to achieve its best average performance. For each test case, the performances are averaged on $200$ Monte–Carlo episodes.
Scenario 1. We consider the following scenario: the transmitter has four antennas (i.e., $M=4$) with unit transmitting power; a communication frame contains $500$ non-pilot data (i.e., the size of the testing data set is $500$); the channel noises are complex Gaussian. First, we suppose that the transmitting antennas emit continuous-valued complex signals; without loss of generality, Gaussian signals are used in experiments. The performance evaluation measure is therefore the mean-squared error (MSE). The experimental results are shown in Fig. 1.
(a) $N$=8, SNR 10dB, $\bm{R}$ Estimated
(b) $N$=8, SNR 10dB, $\bm{R}$ Known
(c) $N$=16, SNR 10dB, $\bm{R}$ Estimated
(d) $N$=16, SNR -10dB, $\bm{R}$ Estimated
Figure 1: Testing MSE against training pilot sizes under different numbers of receiving antennas; only non-robust beamformers including non-diagonally-loaded ones are considered. The true value of $\bm{R}$ can be unknown and estimated using pilot data. The signal-to-noise ratio (SNR) is $10$dB or $-10$dB.
In Fig. 1, for figure clarity, only non-robust beamformers including non-diagonally-loaded ones are considered. From Fig. 1, the following main points can be outlined.
@@
    For a fixed number $M$ of transmitting antennas, the larger the number $N$ of receiving antennas, the smaller the MSE; cf. Figs. 1(a) and 1(c). This fact is well-established and is due to the benefit of antenna diversity. In addition, for fixed $N$ and $M$, the higher the SNR, the smaller the MSE; cf. Figs. 1(c) and 1(d); this is also well believed.
@@
    As the pilot size increases, the Wiener beamformer tends to have the best performance because the Wiener beamformer is optimal for the linear Gaussian signal model. When $\bm{R}$ is accurately known, the Wiener-CE beamformer outperforms the general Wiener beamformer (cf. Fig. 1(b)) because the former also exploits the information of the linear signal model in addition to the pilot data, while the latter only utilizes the pilot data. However, when $\bm{R}$ is estimated using the pilot data, the performances of the general Wiener beamformer and the Wiener-CE beamformer have no significant difference; cf. Figs. 1(a) and 1(c). Therefore, Fig. 1 validates our claim that channel estimation is not a necessary operation in beamforming and estimation of wireless signals; recall Subsection III-A3.
@@
    The ZF beamformer tends to be more efficient as $N$ increases; cf. Figs. 1(a) and 1(c). However, the ZF beamformer becomes less satisfactory when the SNR decreases; cf. Figs. 1(c) and 1(d). The Capon beamformer is also unsatisfactory when $N$ is small or the SNR is low.
@@
Second, we suppose that the transmitting antennas emit discrete-valued symbols from a constellation that is modulated using quadrature phase-shift keying (QPSK). The performance evaluation measure is therefore the symbol error rate (SER). The experimental results are shown in Fig. 2. We find that all the conclusive main points from Fig. 1 can be obtained from Fig. 2 as well: this validates that minimizing MSE reduces SER. In addition, Figs. 2(c) and 2(d) reveal that the Wiener beamformer even slightly works better than the Wiener-CE beamformer when the pilot size is smaller than $15$ because the uncertainty in the estimated $\hat{\bm{R}}$, on the contrary, misleads the latter. Nevertheless, as the pilot size increases, the Wiener-CE beamformer tends to overlap the Wiener beamformer quickly.
(a) $N$=8, SNR 10dB, $\bm{R}$ Estimated
(b) $N$=8, SNR 10dB, $\bm{R}$ Known
(c) $N$=16, SNR 10dB, $\bm{R}$ Estimated
(d) $N$=16, SNR -10dB, $\bm{R}$ Estimated
Figure 2: Testing SER against training pilot sizes under different numbers of receiving antennas; only non-robust beamformers including non-diagonally-loaded ones are considered. The true value of $\bm{R}$ can be unknown and estimated using pilot data. The signal-to-noise ratio (SNR) is $10$dB or $-10$dB.
Scenario 2. Next, we consider another experimental scenario in which impulse channel noises exist. The detailed setups are as follows. The transmitter has four antennas (i.e., $M=4$) with unit transmitting power; without loss of generality, each antenna is assumed to emit continuous-valued complex Gaussian signals. The receiver has eight antennas (i.e., $N=8$). The SNR is $-10$dB, which is a challenging situation; cf. Figs. 2(c) and 2(d). The channel has impulse noises: i.e., in $L$ received signals (i.e., $[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L}]$) that are contaminated by usual complex Gaussian channel noises, $10\%$ of them are also contaminated by uniform noises with the maximum amplitude of $1.5$, which is a relatively large value compared to the amplitude of the usual Gaussian channel noises. As in the first scenario, we assume that a communication frame contains $500$ non-pilot data. To also disclose the computational times of training beamformers, we employ tables to display the experimental results; see Table I and Appendix H (Table II) of the online supplementary materials.
TABLE I: Results of Scenario 2 (Pilot Size = 20)
@@
Wnr 41.42 5.63e-04 Wnr-DL 41.29 2.60e-05
Wnr-DR 38.18 3.51e+00 Wnr-CE 41.42 1.37e-04
Wnr-CE-DL 41.29 4.75e-05 Wnr-CE-DR 41.53 1.33e-04
Capon 86.13 1.15e-04 Capon-DL 85.99 5.23e-05
ZF 75.49 6.89e-05 Kernel 37.20 5.50e-04
Kernel-DL 34.87 1.77e-04
@@
Time: The training time averaged on 200 Monte–Carlo episodes.
Taking the Wiener beamformer for instance, the averaged MSE for each transmitted symbol is $41.42/(500\times 4)=0.021$. From Table I, the following main points can be outlined.
  1. 1.
    Since the signal model under impulse channel noises is no longer linear Gaussian, the optimal beamformer in the MSE sense must be nonlinear. Therefore, the Kernel and the Kernel-DL methods have the potential to outperform other linear beamformers, i.e., to suppress outliers.
  2. 2.
    Distributionally robust beamformers (including diagonally loaded ones) can combat the adverse effect introduced by the limited pilot size and several types of uncertainties in the signal model (e.g., outliers). To be specific, for example, all diagonally-loaded beamformers can outperform their original non-diagonally-loaded counterparts; cf. the Wiener and the Wiener-DL methods, the Wiener-CE and the Wiener-CE-DL methods, the Capon and the Capon-DL methods, and the Kernel and the Kernel-DL methods. In addition, the Wiener-DR beamformer (41) using the nontrivial $F$-norm uncertainty set can further outperform the Wiener-DL beamformer (27) that simply employs the trivial uncertainty set (26).
  3. 3.
    Although the Wiener-DR beamformer has the potential to work better than the Wiener-DL beamformer, it has a significant computational burden, which may not be suitable for timely use in practice especially when the computing resources are limited. Hence, the Wiener-DL beamformer is practically promising (if the pilot size can be improved) because it can provide an excellent balance between the computational burden and the performance.
Remark on Parameter Tuning. From experiments on Scenarios 1 and 2, we find that the uncertainty quantification coefficients $\epsilon$’s (e.g., in diagonal loading) can neither be too large nor too small. When $\epsilon$’s are too large, the beamformers become overly conservative, while when $\epsilon$’s are too small, the beamformers cannot offer sufficient robustness against data scarcity and uncertainties. In both cases of inappropriate $\epsilon$’s, the performances of beamformers degrade significantly. Therefore, $\epsilon$’s must be carefully tuned in practice, and a rigorous method to tune $\epsilon$’s can be the cross-validation method on the training data set (i.e., the pilot data set). If practitioners just pursue satisfaction rather than optimality, the empirical tuning method is recommended to save training time.
@@ VII Conclusions
This paper introduces a unified mathematical framework for beamforming and estimation of wireless signals from the perspective of data-driven machine learning, which reveals that channel estimation is not a necessary operation; the channel-estimation-based beamformers are advantageous only when the pilot size is small and the covariance matrix $\bm{R}$ of the channel noises is (approximately) known. To combat the limited pilot size and several types of uncertainties in the signal model, the distributionally robust (DR) beamforming and estimation framework is then suggested. We prove that the existing diagonal-loading (DL) methods are distributionally robust against the scarcity of pilot data and the uncertainties in the signal model. In addition, we generalize the diagonal-loading methods to achieve better estimation performance (e.g., the DR Wiener beamformer using $F$-norm for uncertainty quantification). Experiments suggest that nonlinear beamformers such as the kernel and the kernel-DL methods have the potential when the pilot size is small and/or the signal model is not linear Gaussian. Compared with the kernel and the kernel-DL beamformers, neural-network-based solutions [10, 12] have the stronger expressive capability of nonlinearities, which however are unscalable in the numbers of transmitting and receiving antennas, and significantly more time-consuming in training and more troublesome in tuning hyper-parameters (e.g., the number of layers and the number of neurons in each layer) than the studied eleven beamformers.

### v2

@@ V-A2 Multi-Frame Case: Dynamic Channel Evolution
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{RKHS-MF}}&=[\bm{R}_{\underline{zs}}+\lambda\bm{W}^{\prime\mathsf{T}}][\bm{R}_{\underline{z}}+\lambda\bm{I}_{L}]^{-1}\\ &=\left(\frac{1}{L}\underline{\bm{S}}\bm{K}+\lambda\bm{W}^{\prime\mathsf{T}}\right)\cdot\left(\frac{1}{L}\bm{K}^{2}+\lambda\bm{I}_{L}\right)^{-1},\end{array}$ (68)
where $\lambda\geq 0$ is a tuning parameter to control the similarity between $\bm{W}$ and $\bm{W}^{\prime}$; cf. Claim 1. $\square$
The remaining distributional robustness modeling and analyses on (67) against the uncertainties in $\hat{\bm{R}}_{\underline{z}}$, $\hat{\bm{R}}_{\underline{xz}}$, and $\hat{\bm{R}}_{\underline{s}}$ are technically straightforward; cf. Subsection IV-C. Therefore, we omit them here.
@@ VI Experiments
The following eleven methods are implemented in the experiments: 1) Wiener: Wiener beamformer (13), upper expression; 2) Wiener-DL: Wiener beamformer with diagonal loading (30), upper expression; 3) Wiener-DR: Distributionally robust Wiener beamformer (49) and (53); 4) Wiener-CE: Channel-estimation-based Wiener beamformer (13), lower expression; 5) Wiener-CE-DL: Channel-estimation-based Wiener beamformer with diagonal loading (30), lower expression; 6) Wiener-CE-DR: Distributionally robust channel-estimation-based Wiener beamformer (41) and (31); 7) Capon: Capon beamformer (38) for $\epsilon_{0}=0$; 8) Capon-DL: Capon beamformer with diagonal loading (38); 9) ZF: Zero-forcing beamformer where $\bm{W}_{\text{ZF}}\coloneqq(\hat{\bm{H}}^{\mathsf{H}}\hat{\bm{H}})^{-1}\hat{\bm{H}}^{\mathsf{H}}$ and $\hat{\bm{H}}$ denotes the estimated channel matrix; 10) Kernel: Kernel receiver (61) with $\epsilon_{0}=0$ in (60); and 11) Kernel-DL: Kernel receiver with diagonal loading (65). Note that the diagonal-loading-based methods are particular cases of distributionally robust beamformers; see, e.g., Corollary 1 and Example 4. The deep-learning-based (DL-based) methods in Subsection V-B are not implemented in this section because they have been deeply studied in our previous publications, e.g., [10, 12]; we only comment on the advantages and disadvantages of DL-based methods compared with the listed eleven methods in Section VII (Conclusions).
@@
When covariance matrix $\bm{R}_{s}$ of transmitted signal $\mathbf{s}$ is unknown for the receiver (e.g., in ISAC systems, $\bm{R}_{s}$ needs to vary from one frame to another for sensing), $\bm{R}_{s}$ is estimated by the sample covariance matrix $\hat{\bm{R}}_{s}=\bm{S}\bm{S}^{\mathsf{H}}/L$. The channel matrix $\bm{H}$ is estimated using the minimum mean-squared error method, i.e., $\hat{\bm{H}}=\bm{X}\bm{S}^{\mathsf{H}}(\bm{S}\bm{S}^{\mathsf{H}})^{-1}$. Covariance matrix $\bm{R}_{v}$ of channel noise $\mathbf{v}$ is estimated using the least-square method, i.e., $\hat{\bm{R}}_{v}=(\bm{X}-\hat{\bm{H}}\bm{S})(\bm{X}-\hat{\bm{H}}\bm{S})^{\mathsf{H}}/L$. The matrices $\hat{\bm{R}}_{s}$, $\hat{\bm{H}}$, and $\hat{\bm{R}}_{v}$ are therefore uncertain compared to their true (but unknown; possibly time-varying) values $\bm{R}_{s}$, $\bm{H}$, and $\bm{R}_{v}$, respectively. The matrices $\hat{\bm{R}}_{s}$, $\hat{\bm{H}}$, and $\hat{\bm{R}}_{v}$ are used in beamformers such as the channel-estimation-based Wiener beamformer (30), the Capon beamformer, and the zero-forcing beamformer.
@@
The beamformers are determined on the training data set (i.e., pilot data). The performance evaluation method of beamformers is mean-squared estimation error (MSE) on the test data set (i.e., non-pilot communication data): to be specific, $\|\bm{S}_{\text{test}}-\hat{\bm{S}}_{\text{test}}\|^{2}_{F}/(M\times L_{\text{test}})$ where $\bm{S}_{\text{test}}\in\mathbb{C}^{M\times L_{\text{test}}}$ is the test data block, $\hat{\bm{S}}_{\text{test}}$ is its estimate, and $L_{\text{test}}$ is the length of non-pilot test data units. As data-driven machine learning methods, all parameters (e.g., uncertainty quantification coefficients $\epsilon$’s) of beamformers can be tuned using the popular cross-validation (e.g., one-shot cross-validation) method. The parameters can also be empirically tuned to save training times because cross-validation imposes a significant computational burden. This article mainly uses the empirical tuning method (i.e., trial-and-error) to tune each beamformer to achieve its best average performance. For each test case, the MSE performances are averaged on $250$ Monte–Carlo episodes.
[…]
@@
    A larger number of pilot data benefits the estimation performances of wireless signals.
@@
    The diagonal loading operation can significantly improve the estimation performances especially when the pilot data size is relatively small.
@@
    Since the signal model under impulse channel noises is no longer linear Gaussian, the optimal beamformer in the MSE sense must be nonlinear. Therefore, the Kernel and the Kernel-DL methods have the potential to outperform other linear beamformers, i.e., to suppress outliers. However, in practice, the non-robust Kernel method may undergo numerical instability in calculating the inverse of the kernel matrix $\bm{K}$. Therefore, its actual MSEs are not necessarily smaller than those of linear beamformers. Nevertheless, the robust Kernel-DL method consistently outperforms all other beamformers.
@@
    Distributionally robust beamformers (including diagonal-loading ones) can combat the adverse effect introduced by the limited pilot size and several types of uncertainties in the signal model (e.g., outliers). To be specific, for example, all diagonal-loading beamformers can outperform their original non-diagonal-loading counterparts; cf. the Wiener and the Wiener-DL methods, the Wiener-CE and the Wiener-CE-DL methods, the Capon and the Capon-DL methods, and the Kernel and the Kernel-DL methods. In addition, the Wiener-DR beamformer (53) using the $F$-norm uncertainty set has the potential to outperform the Wiener-DL beamformer (30) that employs the simple uncertainty set (26).
  5. 5.
    Although the Wiener-DR beamformer has the potential to work better than the Wiener-DL beamformer, it has a significant computational burden, which may not be suitable for timely use in practice especially when the computing resources are limited. Hence, the Wiener-DL beamformer is practically promising because it can provide an excellent balance between the computational burden and the actual performance.
Remarks on Parameter Tuning: From experiments, we find that the uncertainty quantification coefficients $\epsilon$’s (e.g., in diagonal loading) can be neither too large nor too small. When $\epsilon$’s are too large, the beamformers become overly conservative, while when $\epsilon$’s are too small, the beamformers cannot offer sufficient robustness against data scarcity and model uncertainties. In both cases of inappropriate $\epsilon$’s, the performances of beamformers degrade significantly. Therefore, $\epsilon$’s must be carefully tuned in practice, and a rigorous method to tune $\epsilon$’s can be the cross-validation method on the training data set (i.e., the pilot data set). If practitioners just pursue satisfaction rather than optimality, the empirical tuning method is recommended to save training times.
TABLE I: Experimental Results (Pilot Size = 10)
@@
Wnr 3.30 1.49e-04 Wnr-DL 2.11 9.81e-06
Wnr-DR 1.97 3.16e+00 Wnr-CE 3.30 4.59e-05
Wnr-CE-DL 2.50 2.17e-05 Wnr-CE-DR 3.31 4.63e-05
Capon 5.44 4.42e-05 Capon-DL 4.52 2.50e-05
ZF 2.12 2.54e-05 Kernel 1.07 1.60e-04
Kernel-DL 0.80 5.59e-05
@@
Time: The training time averaged on 250 Monte–Carlo episodes.
TABLE II: Experimental Results (Pilot Size = 15)
Beamformer MSE Time Beamformer MSE Time
Wnr 1.38 1.65e-04 Wnr-DL 1.23 1.10e-05
Wnr-DR 1.07 3.21e+00 Wnr-CE 1.38 4.44e-05
Wnr-CE-DL 1.30 2.12e-05 Wnr-CE-DR 1.39 4.28e-05
Capon 4.48 4.31e-05 Capon-DL 4.34 2.42e-05
ZF 2.97 2.44e-05 Kernel 1.12 1.94e-04
Kernel-DL 0.70 9.23e-05
TABLE III: Experimental Results (Pilot Size = 20)
Beamformer MSE Time Beamformer MSE Time
Wnr 1.12 1.86e-04 Wnr-DL 1.05 1.87e-05
Wnr-DR 0.93 7.19e+00 Wnr-CE 1.12 5.78e-05
Wnr-CE-DL 1.08 3.14e-05 Wnr-CE-DR 1.13 6.01e-05
Capon 5.01 5.93e-05 Capon-DL 4.94 3.81e-05
ZF 3.82 3.56e-05 Kernel 1.20 4.48e-04
Kernel-DL 0.66 3.11e-04
TABLE IV: Experimental Results (Pilot Size = 25)
Beamformer MSE Time Beamformer MSE Time
Wnr 0.92 1.41e-04 Wnr-DL 0.88 1.11e-05
Wnr-DR 0.80 4.22e+00 Wnr-CE 0.92 5.02e-05
Wnr-CE-DL 0.90 2.44e-05 Wnr-CE-DR 0.92 4.78e-05
Capon 4.94 4.93e-05 Capon-DL 4.89 2.85e-05
ZF 4.06 2.72e-05 Kernel 1.14 4.26e-04
Kernel-DL 0.60 2.95e-04
TABLE V: Experimental Results (Pilot Size = 50)
Beamformer MSE Time Beamformer MSE Time
Wnr 0.69 1.75e-04 Wnr-DL 0.68 1.85e-05
Wnr-DR 0.65 6.10e+00 Wnr-CE 0.69 5.81e-05
Wnr-CE-DL 0.68 3.03e-05 Wnr-CE-DR 0.70 5.90e-05
Capon 6.95 5.97e-05 Capon-DL 6.93 3.75e-05
ZF 6.36 3.38e-05 Kernel 0.92 1.81e-03
Kernel-DL 0.53 1.67e-03
TABLE VI: Experimental Results (Pilot Size = 100)
Beamformer MSE Time Beamformer MSE Time
Wnr 0.57 3.41e-04 Wnr-DL 0.57 3.64e-05
Wnr-DR 0.55 4.96e+00 Wnr-CE 0.57 6.35e-05
Wnr-CE-DL 0.57 2.93e-05 Wnr-CE-DR 0.58 6.07e-05
Capon 9.89 6.88e-05 Capon-DL 9.88 3.99e-05
ZF 9.45 3.27e-05 Kernel 0.72 5.93e-03
Kernel-DL 0.49 5.83e-03
@@ VII Conclusions
This article introduces a unified mathematical framework for receive beamforming of wireless signals from the perspective of data-driven machine learning, which reveals that channel estimation is not a necessary operation. To combat the limited pilot size and several types of uncertainties in the signal model, the distributionally robust (DR) receive beamforming framework is then suggested. We prove that the diagonal-loading (DL) methods are distributionally robust against the scarcity of pilot data and the uncertainties in the signal model. In addition, we generalize the diagonal-loading methods to achieve better estimation performance (e.g., the DR Wiener beamformer using $F$-norm for uncertainty quantification), at the cost of significantly higher computational burdens. Experiments suggest that nonlinear beamformers such as the Kernel and the Kernel-DL methods have the potential when the pilot size is small and/or the signal model is not linear Gaussian. Compared with the Kernel and the Kernel-DL beamformers, neural-network-based solutions [10, 12] have the stronger expressive capability of nonlinearities, which however are unscalable in the numbers of transmit and receive antennas, and significantly more time-consuming in training and more troublesome in tuning hyper-parameters (e.g., the number of layers and the number of neurons in each layer) than the studied eleven beamformers.